# arXiv:2603.00610v1[cs.SD] 28 Feb 2026

### v1

@@ Abstract
Machine Learning, ICML
@@ 1 Introduction
The rapid advancement of Artificial Intelligence Generated Content (AIGC) has significantly impacted the creative industries, with music generation emerging as one of the cornerstones for numerous commercial applications in music, movie and entertainment industries (Cao et al., 2025; Ren et al., 2025; Ma et al., 2024). Despite the proliferation of sophisticated generative models, evaluating their outputs remains a fundamental challenge. Unlike text or image generation which sometimes focus on functional or purpose, music evaluation is typically aesthetically subjective and complicated, requiring assessments of musicality and adherence to complex, often multimodal, user instructions.
@@
Consequently, evaluation methodologies have struggled to bridge this gap. Traditional metrics like Fréchet Audio Distance (FAD) (Kilgour et al., 2019) operate at the distribution level, failing to provide the sample-level signals necessary for post-training or filtering. More recent approaches, such as SongEval (Yao et al., 2025), PAM (Deshmukh et al., 2024), and various MOS predictors (Jeong et al., 2025), have advanced the field by offering sample-level scoring. However, these efforts remain fragmented and narrowly specialized. They typically focus on isolated attributes (e.g., only caption alignment) and rely on rigid input assumptions, whereas state-of-the-art music generation models already support flexible input combinations, ranging from simple text prompts to interwoven lyrics and audio references, as illustrated in Figure 1, highlighting a growing mismatch between model capabilities and evaluation methodologies.
We argue that effective evaluation requires compositional alignment, defined here not merely as adherence to simultaneous constraints, but as the capability of a unified model to adaptively agree with human preferences across these optional and varying input conditions – human must determine which of two audio samples is superior in terms of musicality or better aligned with the provided compositional instructions when providing ground truth preference. A framework capable of judging this versatility—handling cases where inputs may be text-only, lyric-guided, or audio-referenced—is currently missing. To bridge this gap, our proposed CMI-RewardBench integrates diverse task-specific datasets to vigorously evaluate whether a single reward model can judge generation quality against the heterogeneous instruction sets inherent to modern AIGC flows.
@@
Figure 1: Comparison between traditional music evaluation and the human judge based on compositional multimodal instruction (CMI), which inspire our reward modeling benchmark framework.
@@ 2.1 RLHF for LLMs and MLLMs
Reinforcement Learning from Human Feedback (RLHF) has successfully aligned generative models with human intent by replacing heuristic proxies (e.g., ROUGE) with learned reward models (Stiennon et al., 2020; Ouyang et al., 2022; Bai et al., 2022). This paradigm has successfully extended to multimodal domains, including text-to-image generation (Kirstain et al., 2023; Xu et al., 2023) and video synthesis (Ahn et al., 2024; Liu et al., 2025b), to address aesthetic quality and semantic alignment. Recent research has also applied preference optimization to speech synthesis to improve naturalness and bridge inference-time distribution gaps (Zhang et al., 2024a, 2025b). Concurrently, the use of LLMs as scalable evaluators (Zheng et al., 2023; Gu et al., 2024) has emerged as a robust alternative to human experts. Qhile these works address text, vision, and speech, music evaluation remains under-explored, fragmented. Current music metrics lacking the capacity to judge condition on compositional instructions.
@@ 2.2 Evaluation Metrics of Music
Alignment and LLM Judging. Alignment is primarily measured via contrastive scores like CLAP (Wu et al., 2023), CLaMP3 (Sharma et al., 2024), MuQ-Mulan (Zhu et al., 2025). While music-specific checkpoints improve human-preference alignment (Grötschla et al., 2025), these metrics are largely restricted to text-to-audio pairs and neglect lyrics or audio prompts. Besides, emerging “LLM-as-a-judge” frameworks like AutoMV (Tang et al., 2025) and music recommendation AutoRaters (Chen et al., ) offer scalable, multimodality evaluation with complex instructions. However, these rely on proprietary models and lack an open-source framework for evaluating compositional multiomdal music instructions with lyrics and audio prompts information.
@@ 3.1.2 CMI-Pref and CMI-Pref-Pseudo: Large-Scale Compositional Multimodal Preference Datasets
We propose a new dataset CMI-Pref in our benchmark to capture human preferences when music is conditioned on CMIs including text descriptions, lyrics, and audio prompts, as well as a large-score dataset CMI-Pref-Pseudo with label provided by LLM-judger.
@@
#Duration of audio(hours) 0.83 16.62 140.54 88.30 797.34 133.80
#Duration of reference audio(hours) - - - - 123.95 48.56
@@
Following the annotation protocol described in D, 31 annotators constructed CMI-Pref. We then utilized Qwen3-Omni to pseudo-labels to form CMI-Pref-Pseudo.
@@ 3.3.2 Training Strategy
We train the reward model using a two-stage pipeline that leverages both large-scale pseudo-labeled data and high-quality human annotations. Both MUS and ALI heads are optimized jointly throughout both stages. Given a training sample, we compute the musicality-related loss $\mathcal{L}_{\text{MUS}}$ and the alignment-related loss $\mathcal{L}_{\text{ALI}}$ when applicable. The overall training objective is a combination of the two:
$\mathcal{L}_{\text{total}}=0.5\cdot\mathcal{L}_{\text{MUS}}+0.5\cdot\mathcal{L}_{\text{ALI}}.$ (4)
@@ 4.1.2 Compositional MultiModal Instruction Alignment Evaluation
Detailed Breakdown by Modality. The breakdown of CMI-Pref Subsets highlights the specific alignment skills imparted by our dataset. A critical finding emerges from the model’s performance on fully compositional conditions. When conditioned on all three modalities concurrently (Text + Lyrics + Audio), the model fine-tuned on CMI-Pref achieves an accuracy of 82.40%. This indicates that CMI-Pref provides essential supervision signals for fine-grained multimodal alignment, a capability that generalist models like Gemini 3 Pro (66.8% on similar constraints) currently lack. Furthermore, for pure textual and lyrics adherence tasks (w/o Audio), the variant w/ f.t.: CMI + MusicEval reaches 70.20% accuracy. This suggests that while CMI-Pref establishes a robust baseline for multimodal handling, the inclusion of MusicEval further refines the model’s sensitivity to semantic and lyrical content.
@@ 4.2 The Effectiveness of Different Training Sets
Table 4: Aggregated ablation results with a fixed CMI-RM architecture. Mean SRCC on PAM averages the musicality and text–music alignment regression tasks. Music Arena accuracy uses the combined score $s_{\text{MUS}}{+}s_{\text{ALI}}$ for pairwise prediction. Mean Acc. on CMI-Pref averages musicality and alignment preference accuracy.
@@
Distill only 0.3925 0.5117 64.78% 71.10%
Distill+CMI-Pref 0.6116 0.7315 71.12% 75.90%
@@
Distill+Both 0.5464 0.8266 72.46% 76.05%
Scratch+Both 0.2630 0.4986 71.12% 72.15%
Aggregated metric.
To summarize results across heterogeneous benchmarks, we report one aggregated score per dataset in Table 4. For PAM, we average SRCC over the musicality and text–music alignment regression tasks. For MusicEval, we report SRCC on its musicality MOS. For Music Arena, we report preference accuracy, where the model prediction uses the combined score $s_{\text{MUS}}+s_{\text{ALI}}$ to match the platform’s overall preference signal. For CMI-Pref, we report mean accuracy averaged over musicality and alignment preferences. Detailed per-task numbers are provided in Table 2, Table 3, and Table 4.
@@
(1) CMI-Pref is the primary driver for cross-benchmark generalization. Fine-tuning on CMI-Pref consistently improves all four aggregated metrics: Distill+CMI yields consistent improvements over the Distill baseline on PAM (0.6116 vs. 0.3925), MusicEval (0.7315 vs. 0.5117), Music Arena (71.12% vs. 64.78%), and CMI-Pref (75.90% vs. 71.10%). This suggests that high-quality human preference data, when coupled with compositional conditions, provides highly transferable supervision signals.
(2) MusicEval provides a complementary signal, but is insufficient alone. While Distill+MusicEval alone does not yield consistent gains across all sets, Distill+Both effectively integrates both datasets. This joint training substantially improves MusicEval correlation (to 0.8266) while pushing Music Arena and CMI-Pref to their peak performances (72.46% and 76.05%, respectively). We observe a mild trade-off on PAM compared to Distill+CMI, indicating a slight objective mismatch between PAM and MusicEval that the joint training mechanism partially mitigates, while more diverse data is required for better generalization.
(3) Distillation initialization is universally beneficial. Comparing Distill+Both with Scratch+Both, we observe that distillation initialization yields consistent gains across all benchmarks. Most notably, it improves performance on Music Arena from 71.12% to 72.46%, and significantly boosts both PAM (0.2630 to 0.5464) and CMI-Pref (72.15% to 76.05%). These results suggest that our large-scale pseudo-label pre-training (Distill) establishes a robust prior that not only mitigates overfitting but also transfers effectively to noisy, out-of-distribution scenarios such as Music Arena, thereby establishing Distill+Both as the optimal training strategy.
@@ D.2 Annotation Platform
D.3 Musicality on MusicArena: Decoding User Preference
Method Accuracy (%) AUC MSE
Single Metric Threshold (Threshold = 0)
Musicality Diff 73.4 0.8001 0.1884
Alignment Diff 69.7 0.7714 0.2031
Mus + Ali Diff (Equal) 72.6 0.7962 0.1851
Regression Models (5-fold CV)
Logistic Regression 73.1 0.7991 0.1845
SVM-Linear 72.5 0.7990 0.1846
SVM-RBF 72.6 0.7798 0.1875
Table 15 demonstrates that musicality has an overwhelmingly dominant influence on general preference. A simple thresholding of the musicality difference already yields strong performance, while regression models provide only marginal gains (e.g., +0.3% in accuracy) due to the mitigation of fold-level noise.
@@ Appendix E Discussions on Musicality
E.1 The Definition of Contextual Musicality
Musicality in previous benchmarks, such as SongEval (Yao et al., 2025), isolates absolute acoustic quality. We argue that evaluating ”natural intuitive music preference” requires context.
In our framework, we introduce the concept of Contextual Musicality. Musicality extends beyond mere audio fidelity; it serves as a comprehensive sample of user preference implicitly guided by the generation condition. Creativity, aesthetic alignment, and technical execution collectively shape musicality. By processing the prompt, our model effectively captures the user’s implicit profile and aesthetic appreciation level, making the musicality score a context-aware aesthetic judgment rather than a sterile acoustic measurement.
E.2 Does Prompt Information Contribute to Musicality in CMI-Pref?
To empirically validate whether context improves musicality prediction, we conduct an ablation study using our CMI-Pref dataset. We train two variants of our model on CMI-Pref, optimizing only the musicality head. The first variant drops all prompts during both training and inference (relying purely on self-attention over the evaluated audio), while the second variant utilizes regular conditional training.
Prompt Modality w/ Condition Acc (%) w/o Condition Acc (%) $\Delta$
Text + Audio + Lyrics 82.40 68.80 +13.60
Text + Audio w/o Lyrics 79.20 68.80 +10.40
Text + Lyrics w/o Audio 72.00 76.00 -4.00
Text w/o Audio & Lyrics 68.80 67.20 +1.60
Overall ($N=500$) 75.60 70.20 +5.40
As shown in Table 14, incorporating prompt conditions significantly enhances the overall musicality prediction accuracy (from 70.20% to 75.60%). The improvement is particularly pronounced when reference audio is present. For instance, the “Text + Audio + Lyrics” and “Text + Audio w/o Lyrics” modalities observe substantial gains of +13.60% and +10.40%, respectively. This phenomenon occurs because the reference audio—often a user’s favorite track—acts as a strong “taste anchor,” providing the model with explicit information about the user’s aesthetic standards. Furthermore, detailed text prompts signal a higher level of musical expertise, helping the model penalize specific generation flaws that a sophisticated listener would find unacceptable.
Conversely, we observe an anomalous performance drop ($4.00\%$) in the “Text + Lyrics w/o Audio” setting when conditions are introduced. We hypothesize that this is due to a bottleneck in the text encoder’s ability to process raw lyric structures (e.g., repetitive rhymes and stanza formatting) without the grounding of an acoustic anchor, introducing semantic noise that marginally disrupts the model’s alignment with human aesthetic judgments.
E.3 Musicality on MusicArena

### v2

# CMI-RewardBench: Evaluating Music Reward Models with Compositional Multimodal Instruction

@@ Abstract
music reward modeling, AI music evaluation, compositional multimodal instruction, music generation
@@ 1 Introduction
The rapid advancement of Artificial Intelligence Generated Content (AIGC) has significantly impacted the creative industries, with music generation emerging as one of the cornerstones for numerous commercial applications in music, movie and entertainment industries (Cao et al., 2025; Ren et al., 2025; Ma et al., 2024). Despite the proliferation of sophisticated generative models, evaluating their outputs remains a fundamental challenge. Generally speaking, music evaluation requires assessing both musicality and instruction following. However, advancements in generative models now necessitate this assessment under flexible multimodal conditions, such as text-only, lyric-guided, and audio-referenced inputs.
@@
Consequently, evaluation methodologies have struggled to bridge this gap. Traditional metrics like Fréchet Audio Distance (FAD) (Kilgour et al., 2019) operate at the distribution level, failing to provide the sample-level signals necessary for post-training or filtering. More recent approaches, such as SongEval (Yao et al., 2025), PAM (Deshmukh et al., 2024), and various MOS predictors (Jeong et al., 2025), have advanced the field by offering sample-level scoring. However, these efforts remain fragmented and narrowly specialized. They typically focus on isolated attributes (e.g., only caption alignment) and rely on rigid input assumptions, whereas state-of-the-art music generation models already support flexible input combinations, ranging from simple text prompts to interwoven lyrics and audio references. This growing mismatch between model capabilities and evaluation methodologies is highlighted in Figure 1.
We argue that effective evaluation requires compositional alignment, defined here not merely as adherence to simultaneous constraints, but as the capability of a unified model to adaptively agree with human preferences across these optional and varying input conditions. Specifically, the model should assign scores or rankings that consistently reflect human judgments regarding both musical quality and instruction adherence, regardless of whether inputs are text-only, lyric-guided, or audio-referenced. A framework capable of judging this versatility is currently missing.
To bridge this gap, our proposed CMI-RewardBench integrates diverse task-specific datasets to vigorously evaluate whether a single reward model can judge generation quality against the heterogeneous instruction sets inherent to modern AIGC flows.
@@
Figure 1: Reward models should act as proxies for human evaluation under compositional multimodal instructions (CMI). Human judges evaluate a generated waveform with respect to the provided prompt—text descriptions, lyrics, and/or reference audio—considering both musicality and instruction alignment. However, current reward models are typically fragmented: they either score musicality from audio alone or capture alignment for a single modality pair, and thus cannot reliably proxy human judgments across flexible CMI inputs.
@@ 2.1 RLHF for LLMs and MLLMs
Reinforcement Learning from Human Feedback (RLHF) has successfully aligned generative models with human intent by replacing heuristic proxies (e.g., ROUGE) with learned reward models (Stiennon et al., 2020; Ouyang et al., 2022; Bai et al., 2022). This paradigm has successfully extended to multimodal domains, including text-to-image generation (Kirstain et al., 2023; Xu et al., 2023) and video synthesis (Ahn et al., 2024; Liu et al., 2025b), to address aesthetic quality and semantic alignment. Recent research has also applied preference optimization to speech synthesis to improve naturalness and bridge inference-time distribution gaps (Zhang et al., 2024a, 2025b). Concurrently, the use of LLMs as scalable evaluators (Zheng et al., 2023; Gu et al., 2024) has emerged as a robust alternative to human experts. While these works address text, vision, and speech, music evaluation remains under-explored, fragmented. Current music metrics lacking the capacity to judge condition on compositional instructions.
@@ 2.2 Evaluation Metrics of Music
Alignment and LLM Judging. Alignment is primarily measured via contrastive scores like CLAP (Wu et al., 2023), CLaMP3 (Sharma et al., 2024), MuQ-Mulan (Zhu et al., 2025). While music-specific checkpoints improve human-preference alignment (Grötschla et al., 2025), these metrics are largely restricted to text-to-audio pairs and neglect lyrics or audio prompts. Besides, emerging “LLM-as-a-judge” frameworks like AutoMV (Tang et al., 2025) and music recommendation AutoRaters (Chen et al., ) offer scalable, multimodality evaluation with complex instructions. However, these rely on proprietary models and lack an open-source framework for evaluating compositional multimodal music instructions with lyrics and audio prompts information.
@@ 3.1.2 CMI-Pref and CMI-Pref-Pseudo: Large-Scale Compositional Multimodal Preference Datasets
We propose a new dataset CMI-Pref in our benchmark to capture human preferences when music is conditioned on CMIs including text descriptions, lyrics, and audio prompts, as well as a large-scale dataset CMI-Pref-Pseudo with label provided by LLM-judger.
@@
#Duration of audio(hours) 0.83 16.62 140.54 88.30 808.85 133.80
#Duration of reference audio(hours) - - - - 178.36 48.56
@@
Following the annotation protocol described in Appendix D, 31 annotators constructed CMI-Pref. We then utilized Qwen3-Omni to pseudo-labels to form CMI-Pref-Pseudo.
@@ 3.3.2 Training Strategy
We train the reward model using a two-stage pipeline that leverages both large-scale pseudo-labeled data and high-quality human annotations. Both MUS and ALI heads are optimized jointly throughout both stages. Given a training sample, we compute the musicality-related loss $\mathcal{L}_{\text{MUS}}$ and the alignment-related loss $\mathcal{L}_{\text{ALI}}$ when applicable. The overall training objective is an average of the two:
$\mathcal{L}_{\text{total}}=\frac{1}{2}(\mathcal{L}_{\text{MUS}}+\mathcal{L}_{\text{ALI}})$ (4)
@@ 4.1.2 Compositional MultiModal Instruction Alignment Evaluation
Detailed Breakdown by Modality. The breakdown of CMI-Pref Subsets highlights the specific alignment skills imparted by our dataset. A critical finding emerges from the model’s performance on fully compositional conditions. When conditioned on all three modalities concurrently (Text + Lyrics + Audio), the model fine-tuned on CMI-Pref achieves an accuracy of 82.40%. This indicates that CMI-Pref provides essential supervision signals for fine-grained multimodal alignment, a capability that generalist models like Gemini 3 Pro (66.8% on similar constraints) currently lack. Furthermore, for pure textual and lyrics adherence tasks (w/o Audio), the variant w/ f.t.: CMI + MusicEval reaches 70.20% accuracy. This suggests that while CMI-Pref establishes a robust baseline for multimodal handling, the inclusion of MusicEval further refines the model’s sensitivity to semantic content.
@@ 4.2 The Effectiveness of Different Training Sets
Aggregated metric.
To summarize results across heterogeneous benchmarks, we report one aggregated score per dataset in Table 4. For PAM, we average SRCC over the musicality and text–music alignment regression tasks. For MusicEval, we report SRCC on its musicality MOS. For Music Arena, we report the accuracy of the preference, where the prediction of the model uses the difference between the musicality score $s_{\text{MUS}}$ of the pair. For CMI-Pref, we report mean accuracy averaged over musicality and alignment preferences. Detailed per-task numbers are provided in Table 2, Table 3, and Table 4.
Table 4: Aggregated ablation results with a fixed CMI-RM architecture. Mean SRCC on PAM averages the musicality and text–music alignment regression tasks. Music Arena accuracy uses musicality score $s_{\text{MUS}}$ for pairwise prediction. Mean Acc. on CMI-Pref averages musicality and alignment preference accuracy.
@@
Distill only 0.3925 0.5117 64.25% 71.10%
Distill+CMI-Pref 0.6116 0.7315 71.41% 75.90%
@@
Distill+Both 0.5464 0.8266 73.43% 76.05%
Scratch+Both 0.2630 0.4986 71.34% 72.15%
@@
(1) CMI-Pref is the primary driver for cross-benchmark generalization. Fine-tuning on CMI-Pref consistently improves all four aggregated metrics: Distill+CMI yields consistent improvements over the Distill baseline on PAM (0.6116 vs. 0.3925), MusicEval (0.7315 vs. 0.5117), Music Arena (71.41% vs. 64.25%), and CMI-Pref (75.90% vs. 71.10%). This suggests that high-quality human preference data, when coupled with compositional conditions, provides highly transferable supervision signals.
(2) MusicEval provides a complementary signal, but is insufficient alone. While Distill+MusicEval alone does not yield consistent gains across all sets, Distill+Both effectively integrates both datasets. This joint training substantially improves MusicEval correlation (to 0.8266) while pushing Music Arena and CMI-Pref to their peak performances (73.43% and 76.05%, respectively). We observe a mild trade-off on PAM compared to Distill+CMI, indicating a slight objective mismatch between PAM and MusicEval that the joint training mechanism partially mitigates, while more diverse data is required for better generalization.
(3) Distillation initialization is universally beneficial. Comparing Distill+Both with Scratch+Both, we observe that distillation initialization yields consistent gains across all benchmarks. Most notably, it improves performance on Music Arena from 71.34% to 73.43%, and significantly boosts both PAM (0.2630 to 0.5464) and CMI-Pref (72.15% to 76.05%). These results suggest that our large-scale pseudo-label pre-training (Distill) establishes a robust prior that not only mitigates overfitting but also transfers effectively to noisy, out-of-distribution scenarios such as Music Arena, thereby establishing Distill+Both as the optimal training strategy.
@@ Appendix E Discussions on Musicality
E.1 Musicality on Music Arena
@@
E.2 Musicality: Intrinsic Audio Quality vs. Contextual Preference
In existing datasets and models such as SongEval (Yao et al., 2025), musicality is primarily treated as an absolute measure of “intrinsic audio quality”. These metrics are typically reference-free, focusing exclusively on the audio itself. However, in the context of music generation, musicality often interacts—or even interferes—with the generation context, specifically the multimodal prompts fed into the model.
Compared to passive music listening, when generating music, a user’s perception of whether a piece is “musical” is rarely isolated. Instead, it is inherently influenced by the aesthetic expectations established by their initial creative intent. Therefore, we argue that evaluating musicality in a practical AIGC workflow requires shifting from a context-free measurement to a context-aware judgment.
E.2.1 Empirical Observations on Prompt Contribution
To empirically validate whether prompt information assists in predicting human musicality preferences, we conducted an ablation study using the CMI-Pref dataset. We compared two variants of our model: (1) a variant that drops all prompt conditions during both training and inference (effectively a reference-free evaluator), and (2) a variant utilizing regular conditional training.
As shown in Table 14, our experimental results reveal several key findings:
  • Performance Gain from Context: Incorporating prompt conditions significantly improves the overall musicality prediction accuracy from 70.20% to 75.60% ($\Delta=+5.40\%$). Specifically, the text-only setting also shows a marginal gain (+1.60%).
  • Reference Audio as a Strong Anchor: The improvement is most pronounced when reference audio is present. The “Text + Audio + Lyrics” and “Text + Audio w/o Lyrics” modalities yield substantial gains of +13.60% and +10.40%, respectively. This suggests that reference-based context provides a much clearer signal for aesthetic preference than text alone.
  • Modality Bottleneck in Lyrics: Conversely, the “Text + Lyrics w/o Audio” setting is the only modality that shows a decline (-4.00%). We attribute this to the ineffectiveness of the frozen MuQ-MuLan text encoder, which struggles to process raw lyric structures without acoustic grounding, thereby introducing semantic noise rather than useful context.
  • Comparison with Reference-Free Baselines: As shown in our main results, Table 2, reference-free metrics like SongEval-RM show competitive performance on general acoustic quality but are outperformed by our context-aware model on CMI-Pref. This highlights the limitation of condition-free evaluation in complex compositional tasks.
Prompt Modality w/ Condition Acc (%) w/o Condition Acc (%) $\Delta$
Text + Audio + Lyrics 82.40 68.80 +13.60
Text + Audio w/o Lyrics 79.20 68.80 +10.40
Text + Lyrics w/o Audio 72.00 76.00 -4.00
Text only 68.80 67.20 +1.60
Overall ($N=500$) 75.60 70.20 +5.40
E.2.2 Hypotheses: Why Does Context Assist Musicality Prediction?
We propose two primary hypotheses to explain why the inclusion of prompts—which are in theory independent of absolute audio quality—enhances the model’s predictive performance:
1. Inherent Association in Human Judgment: During annotation, human experts may find it difficult to completely decouple musicality from alignment. For instance, a prompt specifying a “lo-fi aesthetic” might lead an annotator to perceive low-fidelity audio as a musical choice rather than a technical flaw. Thus, the prompt acts as a “taste anchor” that recalibrates the aesthetic scale.
2. Inference Assistance via Non-trivial Shortcuts: From the data correlation perspective, there is an undeniable, intrinsic correlation between the prompt and the evaluated audio, as the latter is directly generated conditioned on the former from a music generation model. Because of this inherent link, the prompt provides critical contextual clues about the expected acoustic features. By attending to the prompt, the model can anticipate the intended musical elements and more easily identify task-specific artifacts that degrade quality but are difficult to detect from the raw audio alone.
E.2.3 Implications for Evaluating Music Creation
Regardless of whether these gains stem from human psychological anchoring or model-side statistical correlation, the results highlight a crucial distinction between music listening and music creation. While standalone acoustic metrics are sufficient for evaluating isolated tracks, the evaluation of AIGC music creation fundamentally requires context. Therefore, multimodal inputs (Text, Lyrics, Audio) should be treated as an integral part of the evaluator’s state. For reward models in these tasks, transitioning from absolute acoustic measurement to context-aware preference is essential for achieving closer alignment with real-world user intent.